OpenStudy (anonymous):

solve the seperable equation: x(dv/dx)= (1-4v^2)/3v

OpenStudy (across):

Given \[x\frac{dv}{dx}=\frac{1-4v^2}{3v}\]Separating variables \[\frac{3v}{1-4v^2}dv=\frac{1}{x}dx\]Integrating both sides \[-\frac{3}{8}log(1-4v^2)=log(x)+c\]Exponentiating both sides \[\frac{1}{(1-4v^2)^{3/8}}=cx\]Solve for v